OpenStudy (anonymous):

what is y=x^2-8x+3 in vertex form???

OpenStudy (anonymous):

Complete the square -2(x^2+4x) - 3 -2(x^2+4x+4-4) - 3 -2(x^2+4x+4) + 8 - 3 -2(x^2+4x+4) + 5 -2(x+2)^2 + 5

jimthompson5910 (jim_thompson5910):

Completing the square is one way to do it. Here is another: y=x^2-8x+3 is in the form y = ax^2 + bx + c where a = 1 b = -8 c = 3 Use the axis of symmetry formula: x = -b/(2a) x = -(-8)/(2*1) x = 4 This is the x coordinate of the vertex. Plug this into y=x^2-8x+3 to find the y coordinate of the vertex. y=x^2-8x+3 y=(4)^2-8(4)+3 y = 16 - 8(4) + 3 y = 16 - 32 + 3 y = -13 This is the y coordinate of the vertex. Therefore, the vertex is (h,k) = (4, -13) which means we know that a = 1 (given) h = 4 (just found this) h = -13 (just found this) Finally you plug these values into the general vertex equation below to get this y = a(x-h)^2 + k y = 1(x-4)^2 + (-13) y = (x-4)^2 - 13

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome